When taking a 10 question multiple choice test. Where each question has 3 possible answers.

yes, it is wrong no, it is not wrong Question 4Points: O out of 1 possible. 1 available on this attempt. This is attempt 1 of 3. When taking a 10 question multiple choice test, where each question has 3 possible answers, it or more questions correct by guessing alone. would be unusual to get Give your answer in the box above as a whole number Submit Question Question 5 Points: 0 out of 1 possible A manufacturer knows that their items have a normally distributed length. with a mean of 13.1 inches and standard deviation of 1.9 inches f one item is chosen at random, what is the probablity that it is less than 9.8 inches long?
Answer #1

4:

Since each question has 3 possible answers so we have

p = 1/3

Here number of correct answer will be binomial distribution with parameters n=10 and p=1/3. The mean and SD is

mu=np=10/3=3.33, sigma=sqrt{np(1-p)}=sqrt{20/9}=1.49

The largest usual value is

μ + 2σ = 3.33 + 2 . 1.49 6.31

Answer : to get 7 or more

5:

The z-score for X = 9.8 is

9.8 -13.1_174 1.9

The required probability is

P(X < 9.8) = P(z <-1.74) = 0.0409